# Copresheaf Topological Neural Networks: A Generalized Deep Learning Framework

### v2

@@ 6 Discussion and Conclusions
We proposed CTNNs, a unified deep learning framework on (un)structured data. By grounding models in copresheaves over CCs, CTNNs generalize GNNs, sheaf networks, and topological models through directional, heterogeneous message passing. Besides theoretical advances, CTNNs offer empirical benefits across diverse tasks laying the principles for multiscale and anisotropic representation learning.
@@
Professor Adrian J. Lew’s contributions to this publication were as a paid consultant and were not part of his Stanford University duties or responsibilities. T. Birdal acknowledges support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. Birdal was supported by a UKRI Future Leaders Fellowship [grant number MR/Y018818/1] as well as a Royal Society Research Grant RG/R1/241402.
@@ Appendix A Notation
This appendix provides a reference summary of the notation and acronyms used throughout the main text. Table 4 details key mathematical symbols, while Table 5 lists abbreviations and their expansions.
@@
Notation Description $\mathcal{S}$ Underlying vertex set of a combinatorial complex (Def. 1) $\mathcal{X}$ Set of nonempty cells $\subseteq\mathcal{P}(\mathcal{S})$ (Def. 1) $\mathcal{P}(\mathcal{S})$ Power set of $\mathcal{S}$ (Def. 1) $\operatorname{rk}:\mathcal{X}\to\mathbb{Z}_{\geq 0}$ Rank function on cells, mapping to non-negative integers (Def. 1) $\mathbb{Z}_{\geq 0}$ Non-negative integers $\mathcal{X}^{k}$$\{x\in\mathcal{X}:\operatorname{rk}(x)=k\}$, the $k$-cells (Def. 1) $\dim\mathcal{X}$ Dimension of the complex, $\max_{x\in\mathcal{X}}\operatorname{rk}(x)$ (Sec. 2.1) $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$ Neighborhood function, mapping cells to sets of neighbor cells (Def. 2) $\mathcal{N}_{\text{adj}}$ Adjacency neighborhood function (Def. 2) $\mathcal{N}_{\text{inc}}$ Incidence neighborhood function (Def. 2) $\mathcal{X}_{\mathcal{N}}$ Effective support of $\mathcal{N}$, $\{x\in\mathcal{X}\mid\mathcal{N}(x)\neq\emptyset\}$ (Sec. 3) $\mathbb{G}\in\{0,1\}^{m\times n}$ Binary neighborhood matrix (Def. 3) $\mathbb{G}^{\mathcal{N}}$ Copresheaf neighborhood matrix, entries $\rho_{z_{i}\to y_{j}}$ or 0 (Def. 7) $\rho_{y\to x}$ Copresheaf morphism $\mathcal{F}(y)\to\mathcal{F}(x)$ for edge $y\to x$ (Def. 4) $\mathbb{A}_{r,k}$ Copresheaf adjacency matrix between $r$-cells (Def. 8) $\mathbf{B}_{r,k}$ Copresheaf incidence matrix between ranks $r$ and $k$ (Def. 8) $C^{k}(\mathcal{X},\mathcal{F})$ $k$-cochain space, $\bigoplus_{x\in\mathcal{X}^{k}}\mathcal{F}(x)$ (Sec. 3) $\operatorname{Hom}(\mathcal{F}(i),\mathcal{F}(j))$ Space of linear maps from $\mathcal{F}(i)$ to $\mathcal{F}(j)$ (Def. 7) $\mathbf{h}_{x}^{(\ell)}$ Feature vector at cell $x$ in layer $\ell$ (Prop. 1) $\alpha$ Learnable message function (Prop. 1) $\beta$ Learnable update function (Prop. 1) $\oplus$ Permutation-invariant aggregator (Prop. 1, Prop. 3) $G=(V,E)$ Directed or undirected graph (Sec. 2.2) $G_{\mathcal{N}}=(\mathcal{X}_{\mathcal{N}},E_{\mathcal{N}})$ Directed graph induced by $\mathcal{N}$, edges $y\to x$ if $y\in\mathcal{N}(x)$ (Def. 6) $\mathcal{F}(x)$, $\mathcal{F}(e)$ Stalks: vector spaces at vertex $x$ or edge $e$ (Def. 4, Def. 5) $\mathcal{F}_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$ Restriction map in a cellular sheaf for $x\leq e$ (Def. 5) $\delta$ Coboundary map, measures local disagreement in cellular sheaf (Sec. 2.2) $\Delta_{\mathcal{F}}$ Sheaf Laplacian, $\Delta_{\mathcal{F}}=\delta^{T}\delta$ (Sec. 2.2) $\mathcal{Y}$, $\mathcal{Z}$ Collections of cells, used in neighborhood matrices (Def. 3, Def. 7) $W_{q}$, $W_{k}$, $W_{v}$ Learnable projection matrices for queries, keys, values in copresheaf self-attention (Prop. 4) $q_{x}$, $k_{x}$, $v_{x}$ Query, key, and value vectors for cell $x$ in copresheaf self-attention (Prop. 4) $a_{xy}$ Attention coefficient for cells $x$ and $y$ in copresheaf self-attention (Prop. 4)
@@ Appendix B Sheaves and Copresheaves on Graphs: A Category Theoretical Look
This appendix provides a category-theoretic exposition of sheaves and copresheaves, emphasizing their definitions within the language of category theory. Additionally, we illustrate the construction of a copresheaf neighborhood matrix through an explicit combinatorial example, instantiating the concepts developed in the main text.
@@ B.1 Copresheaves
Before diving into the technical definition, it’s helpful to think of a copresheaf as a way of assigning data that flows along the structure of a graph—like signals along neurons, or resources in a network. In categorical terms, this structure formalizes the idea of consistently associating elements of some category $\mathcal{C}$.
@@ B.2 Cellular Sheaves
  • to each incidence relation $x\leq e$, a morphism $\rho_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$, called a restriction map,
@@ B.3 Comparison Between Copresheaves and Cellular Sheaves
Copresheaves provide a versatile and powerful framework for machine learning applications across diverse domains, particularly excelling in scenarios where directional data flow and hierarchical dependencies are paramount. Unlike cellular sheaves, which are defined over undirected graphs and enforce consistency through restriction maps $\rho_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$, copresheaves operate on directed graphs, assigning learnable linear maps $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ along edges. This enables anisotropic information propagation, making them ideal for tasks such as physical simulations—where data flows asymmetrically, as in fluid dynamics or heat transfer—or natural language processing, where sequential word dependencies dominate. More importantly, the vertex-centric design, assigning vector spaces $\mathcal{F}(x)$ solely to vertices, aligns seamlessly with message-passing architectures like Graph Neural Networks (GNNs) and Topological Neural Networks (TNNs), allowing these maps to be parameterized and optimized during training. Empirical evidence from our experiments highlights that copresheaf-based models outperform traditional architectures in capturing complex dynamics, demonstrating their superior ability to model spatially varying patterns and long-range dependencies in general applications. See Table 6 for a summary of the comparison between copresheaves and cellular sheaves.
@@
Aspect Copresheaf Cellular sheaf [Hansen and Ghrist, 2019d]
@@ B.4 Copresheaf Neighborhood Matrix Example
Figure 3: (a) A combinatorial complex $\mathcal{X}=(\mathcal{S},\mathcal{X},\operatorname{rk})$ with $\mathcal{S}=\{a,b,c\}$, cells $\mathcal{X}=\{\{a\},\{b\},\{c\},\{a,b\},\{b,c\}\}$. The figure also shows the induced directed graph $G_{\mathcal{N}_{\text{inc}}}=(\mathcal{V}_{\mathcal{N}},E_{\mathcal{N}})$ from the incidence neighborhood structure on the combinatorial complex $\mathcal{X}$. Each arrow $z\to y$ represents a directed edge from a 1-cell to a 0-cell where $y\subset z$, and is associated with a linear map $\rho_{z\to y}$ as part of the copresheaf. (b) The copresheaf neighborhood matrix (CNM) $\mathbb{G}^{\mathcal{N}_{\text{inc}}}$, where rows are indexed by 0-cells $\{a\},\{b\},\{c\}$ and columns by 1-cells $\{a,b\},\{b,c\}$. The matrix entries are linear maps $\rho_{z\to y}$ when $z\in\mathcal{N}_{\text{inc}}(y)$, and 0 otherwise. This matrix supports directional feature propagation from 1-cells to 0-cells.
@@
$E_{\mathcal{N}}=\{y\to x\mid x\in\mathcal{X}_{\mathcal{N}},y\in\mathcal{N}_{\text{inc}}(x)\}=\{\{a,b\}\to\{a\},\{a,b\}\to\{b\},\{b,c\}\to\{b\},\{b,c\}\to\{c\}\}.$
@@
The CNM $\mathbb{G}^{\mathcal{N}}$ for $\mathcal{Y}=\mathcal{X}^{0}$, $\mathcal{Z}=\mathcal{X}^{1}$ is:
$[\mathbb{G}^{\mathcal{N}}]_{i,j}=\begin{cases}\rho_{z_{j}\to y_{i}}&\text{if }z_{j}\in\mathcal{N}_{\text{inc}}(y_{i}),\\ 0&\text{otherwise}.\end{cases}$
@@
$\mathbb{G}^{\mathcal{N}}=\begin{bmatrix}\rho_{\{a,b\}\to\{a\}}&0\\ \rho_{\{a,b\}\to\{b\}}&\rho_{\{b,c\}\to\{b\}}\\ 0&\rho_{\{b,c\}\to\{c\}}\end{bmatrix}.$

### v3

@@ 7 Discussion and Conclusions
We proposed CTNNs, a unified deep learning framework on (un)structured data. By develping models on copresheaves over CCs, CTNNs generalize GNNs, SNNs, and TNNs through directional, heterogeneous message passing. Besides theoretical advances, CTNNs offer empirical benefits across diverse tasks, laying the principles for multiscale and anisotropic representation learning.
@@
Professor Adrian J. Lew’s contributions to this publication were as a paid consultant and were not part of his Stanford University duties or responsibilities. T. Birdal acknowledges support from the Engineering and Physical Sciences Research Council [grant EP/X011364/1]. T. Birdal was supported by a UKRI Future Leaders Fellowship [grant number MR/Y018818/1] as well as a Royal Society Research Grant RG/R1/241402. The authors thank Hans Riess for pointing out the relationship between the CTNNs and the quiver Laplacian.
@@ Appendix A Notation
We provide a reference summary of the notation and acronyms used throughout the main text. Table 5 details key mathematical symbols, while Table 6 lists abbreviations and their expansions.
@@
Notation Description $\mathcal{S}$ Underlying vertex set of a combinatorial complex (Def. 1) $\mathcal{X}$ Set of nonempty cells $\subseteq\mathcal{P}(\mathcal{S})$ (Def. 1) $\mathcal{P}(\mathcal{S})$ Power set of $\mathcal{S}$ (Def. 1) $\operatorname{rk}:\mathcal{X}\to\mathbb{Z}_{\geq 0}$ Rank function on cells, mapping to non-negative integers (Def. 1) $\mathbb{Z}_{\geq 0}$ Non-negative integers $\mathcal{X}^{k}$ $\{x\in\mathcal{X}:\operatorname{rk}(x)=k\}$, the $k$-cells (Def. 1) $\dim\mathcal{X}$ Dimension of the complex, $\max_{x\in\mathcal{X}}\operatorname{rk}(x)$ (Sec. 2.1) $\mathcal{N}:\mathcal{X}\to\mathcal{P}(\mathcal{X})$ Neighborhood function, mapping cells to sets of neighbor cells (Def. 2) $\mathcal{N}_{\text{adj}}$ Adjacency neighborhood function (Def. 2) $\mathcal{N}_{\text{inc}}$ Incidence neighborhood function (Def. 2) $\mathcal{X}_{\mathcal{N}}$ Effective support of $\mathcal{N}$, $\{x\in\mathcal{X}\mid\mathcal{N}(x)\neq\emptyset\}$ (Sec. 3) $\mathbf{G}\in\{0,1\}^{m\times n}$ Binary neighborhood matrix (Def. 3) $\mathbf{G}^{\mathcal{N}}$ Copresheaf neighborhood matrix, entries $\rho_{z_{i}\to y_{j}}$ or 0 (Def. 7) $\rho_{y\to x}$ Copresheaf morphism $\mathcal{F}(y)\to\mathcal{F}(x)$ for edge $y\to x$ (Def. 4) $\mathbf{A}_{r,k}$ Copresheaf adjacency matrix between $r$-cells (Def. 8) $\mathbf{B}_{r,k}$ Copresheaf incidence matrix between ranks $r$ and $k$ (Def. 8) $C^{k}(\mathcal{X},\mathcal{F})$ $k$-cochain space, $\bigoplus_{x\in\mathcal{X}^{k}}\mathcal{F}(x)$ (Sec. 3) $\operatorname{Hom}(\mathcal{F}(i),\mathcal{F}(j))$ Space of linear maps from $\mathcal{F}(i)$ to $\mathcal{F}(j)$ (Def. 7) $\mathbf{h}_{x}^{(\ell)}$ Feature vector at cell $x$ in layer $\ell$ (Prop. 1) $\alpha$ Learnable message function (Prop. 1) $\beta$ Learnable update function (Prop. 1) $\oplus$ Permutation-invariant aggregator (Prop. 1, Prop. 3) $G=(V,E)$ Directed or undirected graph (Sec. 2.2) $G_{\mathcal{N}}=(\mathcal{X}_{\mathcal{N}},E_{\mathcal{N}})$ Directed graph induced by $\mathcal{N}$, edges $y\to x$ if $y\in\mathcal{N}(x)$ (Def. 6) $\mathcal{F}(x)$, $\mathcal{F}(e)$ Stalks: vector spaces at vertex $x$ or edge $e$ (Def. 4, Def. 5) $\mathcal{F}_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$ Restriction map in a cellular sheaf for $x\leq e$ (Def. 5) $\delta$ Coboundary map, measures local disagreement in cellular sheaf (Sec. 2.2) $\Delta_{\mathcal{F}}$ Sheaf Laplacian, $\Delta_{\mathcal{F}}=\delta^{T}\delta$ (Sec. 2.2) $\mathcal{Y}$, $\mathcal{Z}$ Collections of cells, used in neighborhood matrices (Def. 3, Def. 7) $W_{q}$, $W_{k}$, $W_{v}$ Learnable projection matrices for queries, keys, values in copresheaf self-attention (Prop. 4) $q_{x}$, $k_{x}$, $v_{x}$ Query, key, and value vectors for cell $x$ in copresheaf self-attention (Prop. 4) $a_{xy}$ Attention coefficient for cells $x$ and $y$ in copresheaf self-attention (Prop. 4)
@@ Appendix B Sheaves and Copresheaves on Graphs: A Category Theoretical Look
This appendix provides a category-theoretic exposition of sheaves and copresheaves, emphasizing their definitions within the language of category theory. Additionally, we illustrate the construction of copresheaf neighborhood matrices through explicit combinatorial examples, instantiating the concepts developed in the main text.
@@ B.1 Copresheaves
Before diving into the technical definition, it is helpful to think of a copresheaf as a way of assigning data that flows along the structure of a graph, like signals along neurons, or resources in a network. In categorical terms, this structure formalizes the idea of consistently associating elements of some category $\mathcal{C}$.
@@ B.2 Cellular Sheaves
  • to each incidence relation $x\leq e$, a morphism $\mathcal{F}_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$, called a restriction map,
@@ B.3 Comparison Between Copresheaves and Cellular Sheaves
Copresheaves provide a versatile and powerful framework for machine learning applications across diverse domains, particularly excelling in scenarios where directional data flow and hierarchical dependencies are paramount. Unlike cellular sheaves, which are defined over undirected graphs and enforce consistency through restriction maps $\mathcal{F}_{x\leq e}:\mathcal{F}(x)\to\mathcal{F}(e)$, copresheaves operate on directed graphs, assigning learnable linear maps $\rho_{x\to y}:\mathcal{F}(x)\to\mathcal{F}(y)$ along edges. This enables anisotropic information propagation, making them ideal for tasks such as physical simulations, where data flows asymmetrically, as in fluid dynamics or heat transfer, or natural language processing, where sequential word dependencies dominate. More importantly, the vertex-centric design, assigning vector spaces $\mathcal{F}(x)$ solely to vertices, aligns seamlessly with message-passing architectures like Graph Neural Networks (GNNs) and Topological Neural Networks (TNNs), allowing these maps to be parameterized and optimized during training. Empirical evidence from our experiments highlights that copresheaf-based models outperform traditional architectures in capturing complex dynamics, demonstrating their superior ability to model spatially varying patterns and long-range dependencies in general applications. See Table 7 for a summary of the comparison between copresheaves and cellular sheaves.
@@
Aspect Copresheaf Cellular sheaf [Hansen and Ghrist, 2019b]
@@ B.4 Copresheaf Neighborhood Matrix Example
Example 1.
Setup
Let the combinatorial complex $\mathcal{X}=(\mathcal{S},\mathcal{X},\operatorname{rk})$ have symbols $\mathcal{S}=\{a,b,c,d\}$ and cells
$\mathcal{X}^{0}$ $=\big\{\{a\},\{b\},\{c\},\{d\}\big\},$
$\mathcal{X}^{1}$ $=\big\{\{a,b\},\{b,c\},\{c,a\},\{d,b\},\{c,d\}\big\},$
$\mathcal{X}^{2}$ $=\big\{\{a,b,c\},\{d,b,c\}\big\}.$
Ranks satisfy $\operatorname{rk}(\mathcal{X}^{0})=0$, $\operatorname{rk}(\mathcal{X}^{1})=1$, $\operatorname{rk}(\mathcal{X}^{2})=2$. Geometrically, this is the union of two triangles $(a,b,c)$ and $(d,b,c)$ sharing the edge $\{b,c\}$.
Let $\mathcal{E}=[e_{1},e_{2},e_{3},e_{4},e_{5}]$ with
$e_{1}{=}\{a,b\},\qquad e_{2}{=}\{b,c\},\qquad e_{3}{=}\{c,a\},\qquad e_{4}{=}\{d,b\},\qquad e_{5}{=}\{c,d\}.$
Edge–via–face neighborhood
Define $\mathcal{N}_{\triangle}:\mathcal{X}^{1}\to\mathcal{P}(\mathcal{X}^{1})$ by
$\mathcal{N}_{\triangle}(e)\;=\;\Big\{\,e^{\prime}\in\mathcal{X}^{1}:\;\exists\,f\in\mathcal{X}^{2}\text{ with }e\subset f,\;e^{\prime}\subset f,\;e^{\prime}\neq e\,\Big\}.$
Thus, two edges are neighbors iff they both bound the same 2-cell. Concretely,
$\mathcal{N}_{\triangle}(e_{1})$ $=\{e_{2},e_{3}\},\quad\mathcal{N}_{\triangle}(e_{2})=\{e_{1},e_{3},e_{4},e_{5}\},$
$\mathcal{N}_{\triangle}(e_{3})$ $=\{e_{1},e_{2}\},\quad\mathcal{N}_{\triangle}(e_{4})=\{e_{2},e_{5}\},\quad\mathcal{N}_{\triangle}(e_{5})=\{e_{2},e_{4}\}.$
The effective support is $\mathcal{X}_{\mathcal{N}}=\mathcal{X}^{1}$ (edges only).
Induced directed graph
The induced directed graph $G_{\mathcal{N}_{\triangle}}=(V_{\mathcal{N}},E_{\mathcal{N}})$ has vertices $V_{\mathcal{N}}=\mathcal{X}^{1}$ and directed edges
$E_{\mathcal{N}}=\{\,e^{\prime}\to e\;|\;e\in\mathcal{X}^{1},\;e^{\prime}\in\mathcal{N}_{\triangle}(e)\,\}.$
Messages flow from an edge $e^{\prime}$ to an adjacent edge $e$ whenever both bound a common triangle.
Copresheaf on the edge–adjacency poset
Assign to every edge a feature space $\mathcal{F}(e)=\mathbb{R}^{2}$. For each directed adjacency $e^{\prime}\to e$, attach a linear map $\rho_{e^{\prime}\to e}:\mathbb{R}^{2}\to\mathbb{R}^{2}$. To keep notation compact, we write diagonal maps as $\operatorname{diag}(\alpha,\beta)$ and $\mathrm{Id}_{2}$ for the $2\times 2$ identity:
[…]
(Any consistent choice works; the point is a map per directed adjacency $e^{\prime}\to e$.)
Copresheaf Neighborhood Matrix (CNM)
For the ordering $\mathcal{E}=[e_{1},e_{2},e_{3},e_{4},e_{5}]$, the CNM $\mathbf{G}^{\mathcal{N}_{\triangle}}\in\big(\mathbb{R}^{2\times 2}\big)^{5\times 5}$ has block entries
$\big[\mathbf{G}^{\mathcal{N}_{\triangle}}\big]_{i,j}=\begin{cases}\rho_{e_{j}\to e_{i}},&\text{if }e_{j}\in\mathcal{N}_{\triangle}(e_{i}),\\ \mathbf{0}_{2\times 2},&\text{otherwise}.\end{cases}$
Displayed explicity as:
$\mathbf{G}^{\mathcal{N}_{\triangle}}=\left(\begin{array}[]{ccccc}\mathbf{0}_{2\times 2}&\rho_{e_{2}\to e_{1}}&\rho_{e_{3}\to e_{1}}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}\\ \rho_{e_{1}\to e_{2}}&\mathbf{0}_{2\times 2}&\rho_{e_{3}\to e_{2}}&\rho_{e_{4}\to e_{2}}&\rho_{e_{5}\to e_{2}}\\ \rho_{e_{1}\to e_{3}}&\rho_{e_{2}\to e_{3}}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}\\ \mathbf{0}_{2\times 2}&\rho_{e_{2}\to e_{4}}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}&\rho_{e_{5}\to e_{4}}\\ \mathbf{0}_{2\times 2}&\rho_{e_{2}\to e_{5}}&\mathbf{0}_{2\times 2}&\rho_{e_{4}\to e_{5}}&\mathbf{0}_{2\times 2}\end{array}\right)$
See Figure 3.
$e_{1}$$e_{2}$$e_{3}$$e_{4}$$e_{5}$$\scriptstyle a$$\scriptstyle b$$\scriptstyle c$$\scriptstyle d$
$\scriptstyle e_{1}$$\scriptstyle e_{2}$$\scriptstyle e_{3}$$\scriptstyle e_{4}$$\scriptstyle e_{5}$
$\left(\begin{array}[]{ccccc}\mathbf{0}_{2\times 2}&\rho_{e_{2}\to e_{1}}&\rho_{e_{3}\to e_{1}}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}\\ \rho_{e_{1}\to e_{2}}&\mathbf{0}_{2\times 2}&\rho_{e_{3}\to e_{2}}&\rho_{e_{4}\to e_{2}}&\rho_{e_{5}\to e_{2}}\\ \rho_{e_{1}\to e_{3}}&\rho_{e_{2}\to e_{3}}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}\\ \mathbf{0}_{2\times 2}&\rho_{e_{2}\to e_{4}}&\mathbf{0}_{2\times 2}&\mathbf{0}_{2\times 2}&\rho_{e_{5}\to e_{4}}\\ \mathbf{0}_{2\times 2}&\rho_{e_{2}\to e_{5}}&\mathbf{0}_{2\times 2}&\rho_{e_{4}\to e_{5}}&\mathbf{0}_{2\times 2}\end{array}\right)$
Figure 3: (a) A combinatorial complex $\mathcal{X}=(\mathcal{S},\mathcal{X},\operatorname{rk})$ with $\mathcal{S}=\{a,b,c,d\}$, edges $\mathcal{X}^{1}=\{\{a,b\},\{b,c\},\{c,a\},\{d,b\},\{c,d\}\}$, and faces $\mathcal{X}^{2}=\{\{a,b,c\},\{d,b,c\}\}$. (b) Induced edge–adjacency digraph: nodes represent the edges of $\mathcal{X}$ and the edges represent the face adjacencies. (c) Copresheaf neighborhood matrix $\mathbf{G}^{\mathcal{N}_{\triangle}}$.
This CNM performs edge-to-edge directional message passing along face-adjacency: each edge $e_{i}$ aggregates transformed features from its face-adjacent neighbors $e_{j}$ via the maps $\rho_{e_{j}\to e_{i}}$. The shared edge naturally becomes a high-degree conduit between the two triangles.
Example 2.
@@
Figure 4: (a) A combinatorial complex $\mathcal{X}=(\mathcal{S},\mathcal{X},\operatorname{rk})$ with $\mathcal{S}=\{a,b,c\}$, cells $\mathcal{X}=\{\{a\},\{b\},\{c\},\{a,b\},\{b,c\}\}$. The figure also indicates the induced directed graph $G_{\mathcal{N}_{\text{inc}}}=(\mathcal{V}_{\mathcal{N}},E_{\mathcal{N}})$ from the incidence neighborhood structure on the combinatorial complex $\mathcal{X}$. Each arrow $z\to y$ represents a directed edge from a 1-cell to a 0-cell where $y\subset z$, and is associated with a linear map $\rho_{z\to y}$ as part of the copresheaf. (b) The induced directed graph $G_{\mathcal{N}_{\text{inc}}}=(\mathcal{V}_{\mathcal{N}},E_{\mathcal{N}})$ from the incidence neighborhood structure $\mathcal{N}_{\text{inc}}$. (c) The copresheaf neighborhood matrix (CNM) $\mathbf{G}^{\mathcal{N}_{\text{inc}}}$, where rows are indexed by 0-cells $\{a\},\{b\},\{c\}$ and columns by 1-cells $\{a,b\},\{b,c\}$. The matrix entries are linear maps $\rho_{z\to y}$ when $z\in\mathcal{N}_{\text{inc}}(y)$, and 0 otherwise. This matrix supports directional feature propagation from 1-cells to 0-cells.
@@
$E_{\mathcal{N}}$ $=\{y\to x\mid x\in\mathcal{X}_{\mathcal{N}},y\in\mathcal{N}_{\text{inc}}(x)\}$
$=\{\{a,b\}\to\{a\},\{a,b\}\to\{b\},\{b,c\}\to\{b\},\{b,c\}\to\{c\}\}.$
@@
The CNM $\mathbf{G}^{\mathcal{N}}$ for $\mathcal{Y}=\mathcal{X}^{0}$, $\mathcal{Z}=\mathcal{X}^{1}$ is:
[…]
@@
$\mathbf{G}^{\mathcal{N}}=\begin{bmatrix}\rho_{\{a,b\}\to\{a\}}&0\\ \rho_{\{a,b\}\to\{b\}}&\rho_{\{b,c\}\to\{b\}}\\ 0&\rho_{\{b,c\}\to\{c\}}\end{bmatrix}$.